• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name igorko & orchen

    Log entry time 21:29:31 on May 08, 2011

    Entry number 358667

    This entry is a followup to: 358666

    Followups:

    keyword=GUIs for 20uA carbon run

    We compared the HRS, BB and HAND GUIs for the curent run (4315) with old
    runs.
    For BB and HAND we used run 3169 and for the HRSs we used 3397 (4He target).
    
    
    The results for L-HRS, HAND and BB are good and it looks like the system
    recovery is sucssesfull (see plots below).
    NOTE:: we have 5 bars turned off in layer 6 of HAND and also we have the
    "normal" noisy bars on layers 1 and 3.
    
    In the R-HRS we see a shift in the TDC spectra of s1&s2 (bottom plot).
    This could be due to the different target or due to a HW issue. We will
    check it with Vince tomorrow during beam studies.
    


    A copy of this log entry has been emailed to: eip@tauphy.tau.ac.il , sgilad@mit.edu , doug@jlab.org , vasulk@jlab.org , igor.korover@gmail.com , rshneor@jlab.org , israel_yaron@yahoo.com , sbeck@bgu.ac.il, nmuangma



    Figure 1



    Figure 2



    Figure 3



    Figure 4



    Figure 5



    Figure 6



    Figure 7



    Figure 8



    Figure 9